### v2

@@ Abstract
JEL codes: C21, C31
Keywords: Spatial autoregression, varying coefficients, nonparametric inference
1.  Introduction
@@
We then apply our methods in two different settings where varying coefficients have been shown to play an important role. In the first application, we test for constant returns to scale (CRS) in the production function of the Chinese nonmetal mineral industry. This can be seen as a follow-up study to Li et al. (2002), who suggest that their estimate of the returns to scale technology, which varies with managerial expense, is close to being constant (i.e. output elasticities of capital and labor sum to unity) for most observed managerial expense levels. Our analysis confirms their results as we cannot reject the CRS hypothesis using a more general version of their model that permits spatial dependence. This suggests that CRS technology may well be a salient feature of this industry.
@@
The rest of the paper is as follows: Section 2 introduces our method in a baseline regression model allowing for higher-order SAR structure. Section 3 extends the model to allow for spatial error dependence of the Kelejian and Prucha (2007) type. Section 4 allows for nonparametric spatial weights, as in Pinkse et al. (2002), and thereby permits some degree of weight matrix robustness. Section 5 shows how our approach can be adapted to the setting where the spatial lag coefficient varies, like in Malikov and Sun (2017). Section 6 applies our methods to the Chinese nonmetal mineral industry and Boston house prices. Section 7 presents the simulation study, and Section 8 concludes. Proofs are in the appendix, while additional empirical and simulation results are in the online appendix.
2.  Baseline higher-order spatial autoregression
@@ 2.1 Test statistic
We approximate $\delta_{k}(z)$ by $\psi_{k}^{h_{k}}(z)^{\prime}\alpha_{k}^{h_{k}}$, where $\psi_{k}^{h_{k}}(z)=\left(\psi_{k1}(z),\ldots,\psi_{kh_{k}}(z)\right)^{\prime}$ and $\alpha_{k}^{h_{k}}=\left(\alpha_{k1},\ldots,\alpha_{kh_{k}}\right)^{\prime}$, for some basis functions $\psi_{k\ell}(\cdot)$, $\ell=1,\ldots,h_{k}$, $k=1,\ldots,d_{\delta}$. Thus each $\delta_{k}(z)$ is approximated by a linear combination of $h_{k}$ basis functions $\psi_{k}^{h_{k}}$ with coefficients $\alpha_{k}^{h_{k}}$. Define the $d_{\alpha}\times 1$ vector $\psi_{i}\equiv\psi_{i}\left(p_{i},z_{i}\right)=\left(p_{i1}\psi_{1}^{h_{1}}\left(z_{i}\right)^{\prime},\ldots,p_{id_{\delta}}\psi_{d_{\delta}}^{h_{d_{\delta}}}\left(z_{i}\right)^{\prime}\right)^{\prime}$ and $\alpha=\left(\alpha_{1}^{h_{1}}{{}^{\prime}},\ldots,\alpha_{d_{\delta}}^{h_{d_{\delta}}}{{}^{\prime}}\right)^{\prime}$ with $d_{\alpha}=\sum_{k=1}^{d_{\delta}}h_{k}$. Then we can write (2.1) as
@@
where $u_{i}=r_{i}+\epsilon_{i}$ and $r_{i}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$ is the approximation error. Let $X$ and $\Psi$ be matrices with typical rows $x_{i}^{\prime}$ and $\psi_{i}^{\prime}$, and $u$ be with elements $u_{i}$. Writing $Q$ for the $n\times d_{\lambda}$ matrix with typical columns $W_{j}y$, (2.3) can then be written as
@@
where $L=[Q,X,\Psi]$ and $\xi=\left(\lambda^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$. The null to approximate $H_{0}^{true}$ is
@@ 2.2 Asymptotic properties
Assumption 1 controls the approximation error. Sufficient conditions for various cases can be found in Chen (2007). We avoid using specific rates because many types of approximation errors occur in the paper and explicitly allowing different rates of decay introduces extra notation without adding much insight beyond our general conditions. Assumption 2 is fairly standard when establishing a central limit theorem for quadratic forms and is used to check a Lyapunov condition. Assumption 3 imposes moment conditions on regressors and instruments.
@@
3.  SAR with error spatial dependence
@@ 3.1 Test statistic
Assumption 8.
$\mathscr{K}(\cdot):\mathbb{R}\rightarrow[-1,1]$ is a kernel that satisfies $\mathscr{K}(0)=1$, $\mathscr{K}(x)=\mathscr{K}(-x)$, $\mathscr{K}(x)=0$ for $|x|>1$ and $|\mathscr{K}(x)-1|\leq C|x|^{\varrho},|x|\leq 1$, for some $\varrho\geq 1$.
Assumption 7 restricts the spatial dependence in the errors to a manageable degree, see e.g. Kelejian and Prucha (2007) and Delgado and Robinson (2015) for similar assumptions. Assumption 8 is a standard assumption on kernels in the HAC setting, see e.g. Kelejian and Prucha (2007).
@@
Assumption 9.
(a) $\mathrm{E}(\ell^{2})=o\left(n^{2\eta}\right)$ where $\eta<\frac{1}{2}(q-2)/(q-1)$ with $q>4$ in Assumption 2. (b) $\sum_{j=1}^{n}\left|\sigma_{ij}\right|d^{\chi}_{ij,1}<C$ for some $\chi\geq 1$. (c) $d^{*}_{ij,m}=d_{ij,m}+\nu_{ij,m}\geq 0$, with $\left|\nu_{ij,m}\right|<C$ and $\nu_{ij,m}$ independent of $v_{i}$ for all $m=1,\ldots,M$.
@@ 3.2 Asymptotic properties
In all subsequent lemmas/theorems, $q$ is defined in Assumption 2 and $\eta$ is defined in Assumption 9. In Lemma 3.1 and Theorems 3.1-3.2, $d_{\xi}$ is defined as $d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$. Furthermore, for any matrix $A$, let $\|A\|=\left\{\overline{\alpha}\left(A^{\prime}A\right)\right\}^{\frac{1}{2}}$ i.e. the spectral norm of $A$.
@@
Let (3.4) hold. Then, under $H_{0}$, Assumptions 1, 3, 4, 6-9, as $n\rightarrow\infty$,
@@
Under Assumptions 1, 3-10, with
@@ 4.1 Test statistic
Following Pinkse et al. (2002), we approximate $g\left(d_{ij}\right)$ with a series of basis functions
@@
where $g_{in}\equiv g_{i}$ is the $i$-th row of $G$, and, once again dropping $n$-subscripting, accordingly extend (2.3) to
@@
with $u_{i}=r_{iG}+r_{i\delta}+\epsilon_{i}$, where $r_{iG}=\sum_{l=d_{\tau}+1}^{\infty}\tau_{l}\sum_{j\neq i}e_{l}\left(d_{ij}\right)y_{j}$ and $r_{i\delta}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$. In matrix notation, we now have
@@ 4.2 Asymptotic properties
Let (4.7) hold. Then, under $H_{0}$, Assumptions 6-9, 11-13, as $n\rightarrow\infty$,
@@
Define $H_{\ell 3}\equiv H_{\ell 3,n}:\alpha=\alpha^{*}\equiv\nu_{3n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{3n}^{\prime}\Gamma_{3n}\nu_{3n})^{\frac{1}{2}}$, with $\nu_{3n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{3n}$ a $d_{\alpha}\times d_{\alpha}$ matrix defined in the proofs. The following assumption generalizes Assumption 5.
Assumption 14.
@@
Under Assumptions 6-14, with
@@
as $n\rightarrow\infty$, the following hold: (i) Under $H_{0}$, $\mathbb{W}_{3}\overset{d}{\rightarrow}N(0,1)$. (ii) $\mathbb{W}_{3}$ provides a consistent test. (iii) Under the sequence of local alternatives $H_{\ell 3}$, $\mathbb{W}_{3}\overset{d}{\rightarrow}N\left(2^{-1/2},1\right)$.
5.  Varying spatial coefficients
@@ 5.1 Test statistic
where $G(z)=\left[H(z),X,\Psi\right]$, $\gamma=\left(\mu^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$, $\mu=\left(\mu_{1}^{\prime},\ldots,\mu_{d_{\mu}}^{\prime}\right)$, and $u$ has $i$-th element $u_{i}=r_{i\lambda}+r_{i\delta}+\epsilon_{i}$ with $r_{i\lambda}=\sum_{j=1}^{d_{\lambda}}r_{ij\lambda}w_{i,j}^{\prime}y$. Defining $\Lambda_{j}(z)=diag(\lambda_{j}\left(z_{1}\right),\ldots,\\ \lambda_{j}\left(z_{n}\right))$ and $S(z)=I-\sum_{j=1}^{d_{\lambda}}\Lambda_{j}(z)W_{j}$, under the usual invertibility conditions,
@@
where $P_{K_{3}}=K_{3}\left(K_{3}^{\prime}K_{3}\right)^{-1}K_{3}^{\prime}$ with $K_{3}$ being an $n\times J_{3}$ instrument matrix with $J_{3}\geq d_{\gamma}=d_{\mu}+d_{\beta}+d_{\alpha}$, but with $J_{3}$ and $d_{\gamma}$ having the same asymptotic order. Note that we have omitted the $z$ argument for brevity. Let $\mathscr{D}_{4}=\frac{1}{n}R\mathcal{U}_{3}R^{\prime}$ with $\mathcal{U}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\Xi_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$ and $T_{3}=\frac{1}{n}G^{\prime}{P_{K_{3}}}G$, and let $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$ with $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$, where $R=\left[I_{d_{\mu}},0_{d_{\mu}\times\left(d_{\beta}+d_{\alpha}\right)}\right]$ or $R=\left[0_{d_{\alpha}\times\left(d_{\mu}+d_{\beta}\right)},I_{d_{\alpha}}\right]$, and $\hat{\Xi}_{3}$ is the Kelejian and Prucha (2007) estimate of $\Xi_{3}=\frac{1}{n}K_{3}^{\prime}\Sigma K_{3}$, with $\Sigma=E(\epsilon\epsilon^{\prime})$. The test statistic is
@@ 5.2 Asymptotic properties
Let (5.7) hold. Then, under $H_{0}$, Assumptions 6-9, 15-17, as $n\rightarrow\infty$,
@@
The following assumption is the appropriate version of Assumptions 5 and 14 for the model considered in this section.
Assumption 18.
@@
Under Assumptions 6-10 and 15-18, suppose that
@@ 6.1 Testing for the CRS in a production function
Based on the CRS, $H_{0}^{true}:\delta_{1}(z_{i})+\delta_{2}(z_{i})=1$, we reparameterize the model as
@@
$\begin{split}&(\mathbb{W}_{0})\ y_{i}^{*}=\beta_{0}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$
@@
where $\psi_{ik}$ is a basis function, with argument $z_{i}$, for k=1,,h/2k=1,...,h/2. We approximate the unknown functions by polynomial and trigonometric functions of different orders. We define the series approximation based Wald statistic derived from this model as $\mathbb{W}_{0}$. Next, we incorporate the SAR structure and consider:
$y_{i}^{*}=\lambda(z_{i})\sum\limits_{j=1}^{n}w_{ij}y_{j}+\beta_{0}+\delta_{0}(z_{i})+\delta_{2}(z_{i})(p_{i2}-p_{i1})+(\delta_{1}(z_{i})+\delta_{2}(z_{i})-1)p_{i1}+\epsilon_{i}.$
We estimate a number of models and perform tests based on the various configurations and statistics discussed in previous sections:
$\begin{split}&(\mathbb{W}_{1})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{2})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{3})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{4})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$
with $\epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j}$ for $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$. Let $d_{ij,m}^{\ast}$ be the geographical distance between firms $i$ and $j$, and $d_{m}$ represent the 10th percentile of all geographical distances. We use two different row-normalized spatial weight matrices, denoted $W_{n}^{p}$ and $W_{n}^{d}$:
@@
$\begin{split}w_{ij}^{p}=\begin{cases}1\quad\text{if}\ i\ \text{and}\ j\ \text{are in the same city},\\ 0\quad\text{otherwise}.\end{cases}w_{ij}^{d}=\begin{cases}\frac{1}{d_{ij,m}^{*}}\quad\text{if}\ d_{ij,m}^{*}<d_{m},\\ 0\quad\text{otherwise}.\end{cases}\end{split}$
For $\mathbb{W}_{i}$, i=0,,3i=0,...,3, $\alpha_{3}=(\alpha_{31},...\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha_{3}=0$. For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha_{3}=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{h})^{\prime}$. Throughout our empirical analysis, we use $\psi_{ik}(z)=z_{i}^{k}$ as the polynomial function for k=1,,hk=1,...,h, and $\psi_{ik}(z)=[\sin(kz_{i}),\cos(kz_{i})]$ as the trigonometric function k=1,,h/2k=1,...,h/2 for $h=2,4$. To compute $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$, we use the Epanechnikov kernel function. For $\mathbb{W}_{3}$, let $\mathcal{E}_{l}$ be the matrix whose typical entry is, for $i\neq j$ and l=1,,hl=1,...,h, $e_{l}(d_{ij})=[d_{ij,m}^{\ast}]^{l}\mathbf{1}(d_{ij,m}^{\ast}<d_{m})$. Then, $c_{il}=\sum_{j=1}^{n}e_{l}(d_{ij})y_{j}$ for l=1,,hl=1,...,h. For $\mathbb{W}_{4}$, we use $\phi_{il}(z)=\frac{1}{h}[\frac{2}{\pi}\tanh(z_{i})]^{l}$ as the polynomial function and $\phi_{il}(z)=\frac{1}{h}\sin(\frac{z_{i}}{2l})$ as the trigonometric function for l=1,,hl=1,...,h. Then, $h_{il}=y_{i}\sum_{j=1}^{n}w_{ij}\phi_{jl}(z)$. For the IVs, we use $K_{1}=K_{3}=[X,\Psi,W_{n}^{p}X,W_{n}^{d}X]$ with respect to $\mathbb{W}_{1}$, $\mathbb{W}_{2}$ and $\mathbb{W}_{4}$, and $K_{2}=[X,\Psi,\mathcal{E}_{1}X,...,\mathcal{E}_{h}X]$ with respect to $\mathbb{W}_{3}$.

### v3

@@ Abstract
Keywords: Spatial autoregression, varying coefficients, inference
JEL classification: C21, C31
1 Introduction
@@
We then apply our methods in two different settings where varying coefficients have been shown to play an important role. In the first application, we test for constant returns to scale (CRS) in the production function of the Chinese nonmetal mineral industry. This re-visits the empirical study in Li et al. (2002), who showed that output elasticities of capital and labor vary with managerial expense. They also noted that the returns to scale, which is the sum of output elasticities of inputs, are close to one at many managerial expense levels suggesting CRS, but they did not have a test for this hypothesis. Using a newer dataset, our methods can test for CRS while allowing for spatial dependence. We do not reject the CRS hypothesis in all settings, which suggests that CRS technology is a salient feature of this industry.
@@
The rest of the paper is as follows: Section 2 introduces our method in a baseline regression model allowing for higher-order SAR structure. Section 3 extends the model to allow for spatial error dependence of the Kelejian and Prucha (2007) type. Section 4 allows for nonparametric spatial weights, as in Pinkse et al. (2002), and thereby permits some degree of weight matrix robustness. Section 5 shows how our approach can be adapted to the setting where the spatial lag coefficient varies, like in Malikov and Sun (2017). Section 6 applies our methods to the Chinese nonmetal mineral industry and Boston house prices. Section 7 presents the simulation study, and Section 8 concludes. Proofs and additional simulation results are in the online appendix.
2 Baseline higher-order spatial autoregression
@@ 2.1 Test statistic
We approximate $\delta_{k}(z)$ by $\psi_{k}^{h_{k}}(z)^{\prime}\alpha_{k}^{h_{k}}$, where $\psi_{k}^{h_{k}}(z)=\left(\psi_{k1}(z),\ldots,\psi_{kh_{k}}(z)\right)^{\prime}$ and $\alpha_{k}^{h_{k}}=\left(\alpha_{k1},\ldots,\alpha_{kh_{k}}\right)^{\prime}$, for some basis functions $\psi_{k\ell}(\cdot)$, $\ell=1,\ldots,h_{k}$, $k=1,\ldots,d_{\delta}$. Thus each $\delta_{k}(z)$ is approximated by a linear combination of $h_{k}$ basis functions $\psi_{k}^{h_{k}}$ with coefficients $\alpha_{k}^{h_{k}}$. Define the $d_{\alpha}\times 1$ vector $\psi_{i}\equiv\psi_{i}\left(p_{i},z_{i}\right)=\left(p_{i1}\psi_{1}^{h_{1}}\left(z_{i}\right)^{\prime},\ldots,p_{id_{\delta}}\psi_{d_{\delta}}^{h_{d_{\delta}}}\left(z_{i}\right)^{\prime}\right)^{\prime}$ and $\alpha=\left(\alpha_{1}^{h_{1}}{{}^{\prime}},\ldots,\alpha_{d_{\delta}}^{h_{d_{\delta}}}{{}^{\prime}}\right)^{\prime}$, with $d_{\alpha}=\sum_{k=1}^{d_{\delta}}h_{k}$. Then we can write (2.1) as
@@
with $u_{i}=r_{i}+\epsilon_{i}$, where $r_{i}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$ is the approximation error. Let $X$ and $\Psi$ be matrices with typical rows $x_{i}^{\prime}$ and $\psi_{i}^{\prime}$, and $u$ be with elements $u_{i}$. Writing $Q$ for the $n\times d_{\lambda}$ matrix with typical columns $W_{j}y$, (2.3) can be written in matrix notation as
@@
where $L=[Q,X,\Psi]$ and $\xi=\left(\lambda^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$. The null hypothesis to approximate $H_{0}^{true}$ is
@@ 2.2 Asymptotic properties
Assumption 2.1 controls the approximation error. Sufficient conditions for various cases can be found in Chen (2007). We avoid using specific rates because many types of approximation errors occur in the paper and explicitly allowing different rates of decay introduces extra notation without adding much insight beyond our general conditions. Assumption 2.2 is fairly standard when establishing a central limit theorem for quadratic forms and is used to check a Lyapunov condition. Assumption 2.3 imposes moment conditions on regressors and instruments.
@@
3 SAR with error spatial dependence
@@ 3.1 Test statistic
Assumption 3.3.
$\mathscr{K}(\cdot):\mathbb{R}\rightarrow[-1,1]$ is a kernel function that satisfies $\mathscr{K}(0)=1$, $\mathscr{K}(x)=\mathscr{K}(-x)$, $\mathscr{K}(x)=0$ for $|x|>1$ and $|\mathscr{K}(x)-1|\leq C|x|^{\varrho},|x|\leq 1$, for some $\varrho\geq 1$.
Assumption 3.2 restricts the spatial dependence in the errors to a manageable degree, see e.g. Kelejian and Prucha (2007) and Delgado and Robinson (2015) for similar assumptions. Assumption 3.3 is a standard assumption on kernels in the HAC setting, see e.g. Kelejian and Prucha (2007).
@@
Assumption 3.4.
(a) $\mathrm{E}(\ell^{2})=o\left(n^{2\eta}\right)$ where $\eta<\frac{1}{2}(q-2)/(q-1)$ with $q>4$ in Assumption 2.2. (b) $\sum_{j=1}^{n}\left|\sigma_{ij}\right|d^{\chi}_{ij,1}<C$ for some $\chi\geq 1$. (c) $d^{*}_{ij,m}=d_{ij,m}+\nu_{ij,m}\geq 0$, with $\left|\nu_{ij,m}\right|<C$ and $\nu_{ij,m}$ independent of $v_{i}$ for all $m=1,\ldots,M$.
@@ 3.2 Asymptotic properties
In all subsequent lemmas/theorems, $q$ is defined in Assumption 2.2 and $\eta$ is defined in Assumption 3.4. In Lemma 3.1 and Theorems 3.1-3.2, $d_{\xi}$ is defined as $d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$. Furthermore, for any matrix $A$, let $\|A\|=\left\{\overline{\alpha}\left(A^{\prime}A\right)\right\}^{\frac{1}{2}}$ i.e. the spectral norm of $A$.
@@
Let (3.4) hold. Then, under $H_{0}$, Assumptions 2.1, 2.3, 2.4, 3.1-3.4, as $n\rightarrow\infty$,
@@
Under Assumptions 2.1, 2.3-3.5, with
@@ 4.1 Test statistic
Following Pinkse et al. (2002), we approximate $g\left(d_{ij}\right)$ of $G$ with a series of basis functions
@@
and (2.3) to
@@
with $u_{i}=r_{iG}+r_{i\delta}+\epsilon_{i}$, where $r_{iG}=\sum_{l=d_{\tau}+1}^{\infty}\tau_{l}\sum_{j\neq i}e_{l}\left(d_{ij}\right)y_{j}$ and $r_{i\delta}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$. In the matrix notation, we now have
@@ 4.2 Asymptotic properties
Let (4.7) hold. Then, under $H_{0}$, Assumptions 3.1-3.4, 4.1-4.3, as $n\rightarrow\infty$,
@@
Define $H_{\ell 3}\equiv H_{\ell 3,n}:\alpha=\alpha^{*}\equiv\nu_{3n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{3n}^{\prime}\Gamma_{3n}\nu_{3n})^{\frac{1}{2}}$, with $\nu_{3n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{3n}$ a $d_{\alpha}\times d_{\alpha}$ matrix defined in the proofs. The following assumption generalizes Assumption 2.5.
Assumption 4.4.
@@
Under Assumptions 3.1-4.4, with
@@
as $n\rightarrow\infty$, the following hold: (1) Under $H_{0}$, $\mathbb{W}_{3}\overset{d}{\rightarrow}N(0,1)$. (2) $\mathbb{W}_{3}$ provides a consistent test. (3) Under the sequence of local alternatives $H_{\ell 3}$, $\mathbb{W}_{3}\overset{d}{\rightarrow}N\left(2^{-1/2},1\right)$.
5 Varying spatial coefficients
@@ 5.1 Test statistic
where $G(z)=\left[H(z),X,\Psi\right]$, $\gamma=\left(\mu^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$, $\mu=\left(\mu_{1}^{\prime},\ldots,\mu_{d_{\mu}}^{\prime}\right)$, and $u$ has $i$-th element $u_{i}=r_{i\lambda}+r_{i\delta}+\epsilon_{i}$ with $r_{i\lambda}=\sum_{j=1}^{d_{\lambda}}r_{ij\lambda}w_{i,j}^{\prime}y$. Defining $\Lambda_{j}(z)=diag\left(\lambda_{j}\left(z_{1}\right),\ldots,\lambda_{j}\left(z_{n}\right)\right)$ and $S(z)=I-\sum_{j=1}^{d_{\lambda}}\Lambda_{j}(z)W_{j}$, under the usual invertibility conditions,
@@
where $P_{K_{3}}=K_{3}\left(K_{3}^{\prime}K_{3}\right)^{-1}K_{3}^{\prime}$ with $K_{3}$ being an $n\times J_{3}$ instrument matrix with $J_{3}\geq d_{\gamma}=d_{\mu}+d_{\beta}+d_{\alpha}$, but with $J_{3}$ and $d_{\gamma}$ having the same asymptotic order. Note that we have omitted the $z$ argument for brevity. Let $\mathscr{D}_{4}=\frac{1}{n}R\mathcal{U}_{3}R^{\prime}$ with $\mathcal{U}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\Xi_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$ and $T_{3}=\frac{1}{n}G^{\prime}{P_{K_{3}}}G$, and let $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$ with $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$, where $R=\left[I_{d_{\mu}},0_{d_{\mu}\times\left(d_{\beta}+d_{\alpha}\right)}\right]$ or $R=\left[0_{d_{\alpha}\times\left(d_{\mu}+d_{\beta}\right)},I_{d_{\alpha}}\right]$, and $\hat{\Xi}_{3}$ is the Kelejian and Prucha (2007) estimate of $\Xi_{3}=\frac{1}{n}K_{3}^{\prime}\Sigma K_{3}$, with $\Sigma=E(\epsilon\epsilon^{\prime})$. Then, the test statistic is
@@ 5.2 Asymptotic properties
Let (5.7) hold. Then, under $H_{0}$, Assumptions 3.1-3.4, 5.1-5.3, as $n\rightarrow\infty$,
@@
The following assumption is the appropriate version of Assumptions 2.5 and 4.4 for the model considered in this section.
Assumption 5.4.
@@
Under Assumptions 3.1-3.5 and 5.1-5.4, suppose that
@@ 6.1 Testing for the CRS in a production function
Based on the CRS hypothesis, $H_{0}^{true}:\delta_{1}(z_{i})+\delta_{2}(z_{i})=1$, we reparameterize the model as
@@
$\begin{split}&y_{i}^{*}=\beta_{0}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$
@@
where $\psi_{ik}$ is a basis function, with argument $z_{i}$, for k=1,,h/2k=1,...,h/2. We approximate the unknown functions by polynomial and trigonometric functions of different orders. We define the series approximation based Wald statistic derived from this model as $\mathbb{W}_{0}$. Next, we estimate a number of models and perform tests based on the various configurations and statistics discussed in previous sections:
$y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i}.$ (6.2)
We compute the Wald statistics as follows: $\mathbb{W}_{1}$ from (6.2) with $i.i.d$ errors $\epsilon_{i}$; $\mathbb{W}_{2}$ from (6.2) with $\epsilon_{i}=\sum_{j=1}^{n}b_{ij}v_{j};$ $\mathbb{W}_{3}$ from (6.2) replacing $\lambda\sum_{j=1}^{n}w_{ij}y_{j}$ by $\sum_{l=1}^{d_{\tau}}\tau_{l}c_{il}$ and setting $\epsilon_{i}=\sum_{j=1}^{n}b_{ij}v_{j};$ $\mathbb{W}_{4}$ from (6.2) with $\lambda\sum_{j=1}^{n}w_{ij}y_{j}$ replaced by $\sum_{m=1}^{d_{\mu}}\mu_{m}\mathfrak{H}_{im}$ and $\varepsilon_{i}=\sum_{j=1}^{n}b_{ij}v_{j}$, where $\mathfrak{H}_{im}$ is the $(i,m)$-th entry of the $n\times d_{\mu}$ matrix $H(z)$ defined in Section 5. The coefficient on $p_{i1}$ remains the measure of returns to scale technology even with a SAR feature, although this notion differs from the overall returns to scale that combines technological and other spillovers effects. Let $d_{ij,m}^{\ast}$ be the geographical distance between firms $i$ and $j$, and $d_{m}$ represent the 10th percentile of all geographical distances. We use two different row-normalized spatial weight matrices, denoted
@@
$\begin{split}w_{ij}^{p}=[W_{n}^{p}]_{ij}=\begin{cases}1\quad\text{if}\ i\ \text{and}\ j\ \text{ in same city},\\ 0\quad\text{otherwise}.\end{cases}w_{ij}^{d}=[W_{n}^{d}]_{ij}=\begin{cases}\frac{1}{d_{ij,m}^{*}}\quad\text{if}\ d_{ij,m}^{*}<d_{m},\\ 0\quad\text{otherwise}.\end{cases}\end{split}$